Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used in production (not on the basis of raw materials purchased). Its predetermined overhead rate was based on a cost formula that estimated $128,800 of manufacturing overhead for an estimated allocation base of $92,000 direct material dollars to be used in production. The company has provided the following data for the just completed year: Purchase of raw materials $ 133,000 Direct labor cost $ 88,000 Manufacturing overhead costs: Indirect labor $ 122,100 Property taxes $ 8,400 Depreciation of equipment $ 20,000 Maintenance $ 16,000 Insurance $ 12,000 Rent, building $ 32,000 Beginning Ending Raw Materials $ 29,000 $ 14,000 Work in Process $ 47,000 $ 38,000 Finished Goods $ 73,000 $ 59,000 Required: 1. Compute the predetermined overhead rate for the year. 2. Compute the amount of underapplied or overapplied overhead for the year. 3. Prepare a schedule of cost of goods manufactured for the year. Assume all raw materials are used in production as direct materials. 4. Compute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answer. 5. Assume that the $38,000 ending balance in Work in Process includes $8,300 of direct materials. Given this assumption, supply the information missing below:

Prepare a schedule of cost of goods manufactured for the year. Assume all raw materials are used in production as direct materials.

Gitano Products
Schedule of Cost of Goods Manufactured
Direct materials:
Total raw materials available
Raw materials used in production
Total manufacturing costs
Cost of goods manufactured

  

Compute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answer.

Unadjusted cost of goods sold

Assume that the $38,000 ending balance in Work in Process includes $8,300 of direct materials. Given this assumption, supply the information missing below:

Direct materials
Direct labor
Manufacturing overhead
Work in process inventory $0
Answer #1
1
Estimated manufacturing overhead 128800
Divide by estimated direct materials 92000
Predetermined overhead rate 140%
2
Indirect labor 122100
Property taxes 8400
Depreciation of equipment 20000
Maintenance 16000
Insurance 12000
Rent, building 32000
Total Manufacturing overhead costs 210500
Total Actual Manufacturing overhead costs 210500
Less: Manufacturing overhead applied to work in process 207200 =148000*140%
Underapplied overhead 3300
3
Schedule of Cost of Goods Manufactured
Direct materials:
Beginning Raw materials inventory 29000
Add: Purchases of raw materials 133000
Total raw materials available 162000
Less: Ending Raw materials inventory 14000
Raw materials used in production 148000
Direct labor cost 88000
Manufacturing overhead applied to work in process 207200 =148000*140%
Total manufacturing costs 443200
Add: Beginning Work in process inventory 47000
490200
Less: Ending Work in process inventory 38000
Cost of goods manufactured 452200
4
Finished goods inventory, beginning 73000
Add: Cost of goods manufactured 452200
Cost of Goods available for sale 525200
Less: Finished goods inventory, ending 59000
Unadjusted Cost of goods sold 466200
5
Direct materials 8300
Direct labor 18080 =38000-8300-11620
Manufacturing overhead 11620 =8300*140%
Work in process inventory 38000